Using the method of variation of parameters, solve the equation:


d²y/dx² +a²y = sec ax.

1
Expert's answer
2022-01-26T17:03:52-0500

"\text{ We have }\\
\left(D^{2}+a^{2}\right) y=\sec a x\\
\text{A.E. is } m^{2}+a^{2}=0 \Rightarrow m=\pm ai\\
\text{C.F. } =y_{c}=C_{1} \cos a x+C 2 \sin a x\\
y=A \cos a x+B \sin a x \ldots(1)\\
\text{ be the complete solution of the given equation where } A \text{ and } B \text{ are to be found.}\\
\text{We have, } y_{1}=\cos a x, y_{2}=\sin a x\\
y_{1}^{\prime}=-a \sin a x, y_{2}^{\prime}=a \cos a x\\
\begin{aligned}
&W=y_{1} y_{2}^{\prime}-y_{2} y_{1}^{\prime}=a . \text { Also, } \phi(x)=\sec a x \\
&A^{\prime}=\frac{-y_{2} \phi(x)}{W}, \text { and } B^{\prime}=\frac{y_{1} \phi(x)}{W} \\
&A^{\prime}=\frac{-\sin a x \cdot \sec a x}{a}, \quad B^{\prime}=\frac{\cos a x \cdot \sec a x}{a}
\end{aligned}
\\[4mm]
\begin{aligned}
A^{\prime} &=\frac{-\tan a x}{a}, \quad B^{\prime}=\frac{1}{a} \\
B &=\frac{1}{a} \int d x+c_{2} \\
A &=-\frac{1}{a} \int \tan a x d x+C_{1}, \quad B=\frac{x}{a}+C_{2} \\
A &=\frac{-\log (\sec a x)}{a^{2}}+C_{1}
\end{aligned}
\\[4mm]
\text{Substituting these values of } A \text{ and } B \text{ in Eqn. (1), we get: }\\"



"\begin{aligned}
y=& C_{1} \cos a x+C_{2} \sin a x 
&-\frac{\cos a x \log (\sec a x)}{a^{2}}+\frac{x \sin a x}{a}
\end{aligned}"